# Constrained mean-field control with singular control: Existence, stochastic maximum principle and constrained FBSDE

### v2

@@ 4 SMP by the Lagrange Multipliers Method
$p_{t}$ $=-\sum_{i\in I}\int_{t}^{T}\left(\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})+\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s}){\eta^{i}_{s}}^{\prime}\right]\right)\mathrm{d}s$
@@
where $MN:=\sum_{i=1}^{r}M_{i}L_{i}\in\mathbb{R}^{n}$ for any $M=(M_{1},\ldots,M_{r})\in\mathbb{R}^{n\times n\times r}$ and $L=(L_{1},\ldots,L_{r})\in\mathbb{R}^{n\times r}$. By construction of the above $L\in\mathbb{T}^{2}$, we have
$\langle\lambda_{1}^{1}(t),L_{t}^{1}\rangle\mathrm{d}t=\langle\lambda_{1}^{1}(t)-p(t),\mathrm{d}L_{t}\rangle+\langle p(t),\mathrm{d}L_{t}\rangle-\langle\lambda_{1}^{1},\nu\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathcal{M}[0,T])^{n}}.$ (24)
@@
$\mathbb{E}\left[\int_{0}^{T}\langle p_{t},\mathrm{d}L_{t}\rangle\right]=-\mathbb{E}\left[\int_{0}^{T}\langle L_{t},\mathrm{d}p_{t}\rangle\mathrm{d}t\right]-\mathbb{E}\left[\langle p,L\rangle_{T}\right]$
@@
$\mathbb{E}\left[\langle p_{0}-\lambda_{1}(0),L_{0}\rangle+\int_{0}^{T}\langle p_{t}-\lambda_{1}^{1}(t),\mathrm{d}L_{t}\rangle+\langle\lambda_{2}-\lambda_{1}^{1},\nu\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathcal{M}[0,T])^{n}}\right]=0.$
We thus deduce that $p_{0}=\lambda_{1}(0)$ and $p$ is indistinguishable from $\lambda_{1}^{1}$ by the arbitrariness of $L$. As a result, $\lambda_{1}^{1}$ solves (5.3). Moreover, we have $G^{\top}\lambda_{2}=G^{\top}\lambda_{1}^{1}$ in $(C([0,T];\mathbb{R})^{**})^{k}$, and hence $G^{\top}\lambda_{2}\in C([0,T];\mathbb{R}^{k})$. Here, $G$ can be viewed a mapping in $\mathcal{L}((C([0,T];\mathbb{R})^{**})^{n};$ $((C([0,T];\mathbb{R})^{**})^{k}))$. By the optimality condition for $\widehat{\alpha}\in\mathscr{U}[0,T]$, we have
$0$ $\leq\mathbb{E}\left[\int_{0}^{T}\left\langle r_{0}\nabla_{u}f(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})+\nabla_{u}b(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})^{\top}\lambda_{t}^{1}+\nabla_{u}\sigma(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})^{\top}\lambda_{t}^{2},\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t\right$.
@@
$0$ $\leq-\mathbb{E}\left[\langle\xi_{3},\eta-\widehat{\zeta}\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathbb{M}[0,T])^{n}}\right]$
$=\mathbb{E}\left[r_{0}\int_{0}^{T}c(t)\mathrm{d}(\eta_{t}-\hat{\zeta}_{t})+\left\langle\lambda_{2},\int_{0}^{T}G(t)\mathrm{d}(\eta_{t}-\hat{\zeta}_{t})\right\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathbb{M}[0,T])^{n}}\right]$
@@ 5 Uniqueness and Stability of Solution to Constrained FBSDE
$Y_{t}$ $=\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\mu},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H^{r_{0}}(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime},Y_{s}^{\prime},Z_{s}^{\prime})(\widehat{X}_{s})\right]\mathrm{d}s$
@@
subject to $|r_{0}|+\|\eta\|_{\mathbb{L}_{\mathbb{F},\mathbb{R}}}=1$, $\widehat{\mu}_{t}=\mathcal{L}(\widehat{X}_{t})$, $\mathbb{P}\left(G(t)Y_{t}+c(t)\geq 0,~\forall t\in[0,T]\right)=1$, and $\mathrm{d}t\otimes\mathrm{d}\mathbb{P}$-a.s.
@@
We now verify the CQ condition as in (1.4) of [33] to establish the KKT optimality condition (i.e., we verify $r_{0}=1$ in Lemma 4.1).
@@
Here, the Hamiltonian $H(t,x,\mu,u,y,z)=H^{1}(t,x,\mu,u,y,z)$ for $(t,x,\mu,u,y,z)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$, and the process pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ is the unique solution to the following general type of constrained BSDE:
$Y_{t}$ $=\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime},Y_{s}^{\prime},Z_{s}^{\prime})(\widehat{X}_{s})\right]\mathrm{d}s$
@@
The strict convexity imposed on the Hamiltonian $H$ w.r.t. the control variable $u\in U$ given in (C3) of Assumption 3 yields the result that once $\langle\nabla_{u}H(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t},Y_{t},Z_{t}),u-\widehat{\alpha}_{t}\rangle\geq 0$ for all $u\in U$, we must have $\widehat{\alpha}_{t}=\hat{u}(t,\widehat{X}_{t},\widehat{\mu}_{t},Y_{t},Z_{t})$. Then, we have
@@
Let $(X^{i}_{t},\zeta^{i}_{t},Y^{i}_{t},Z^{i}_{t},\eta_{t}^{i})_{t\in[0,T]}$ ($i=1,2$) be two solutions to the FBSDE (5)-(27) satisfying $X^{1}_{0}=X_{0}^{2}$ and $Y_{T}^{1}=Y_{T}^{2}$. Set $\bar{X}_{t}:=X_{t}^{1}-X_{t}^{2}$, $\bar{\zeta}_{t}:=\zeta_{t}^{1}-\zeta_{t}^{2}$, $\bar{Y}_{t}:=Y_{t}^{1}-Y_{t}^{2}$, $\bar{Z}_{t}:=Z_{t}^{1}-Z_{t}^{2}$ and $\bar{\eta}_{t}:=\eta_{t}^{1}-\eta_{t}^{2}$ for $t\in[0,T]$. Applying Itô’s formula to $\bar{X}_{t}\bar{Y}_{t}$, we arrive at
$\bar{X}_{T}\bar{Y}_{T}$ $=\int_{0}^{T}\left\{-\Delta\bar{F}(t)\bar{X}_{t}+\bar{Y}\Delta\bar{b}(t)+\mathrm{tr}(\bar{Z}_{t}\Delta\bar{\sigma}(t))\right\}\mathrm{d}t+\int_{0}^{T}\left(\bar{X}_{t}\bar{Z}_{t}+\bar{Y}_{t}\Delta\sigma(t)\right)\mathrm{d}W_{t}$
@@
where $\Delta\bar{F}(t):=\bar{F}(t,X_{t}^{1},\mathcal{L}(X_{t}^{1}),Y_{t}^{1},Z_{t}^{1})-\bar{F}(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})$, $\Delta\bar{b}(t):=\bar{b}(t,X_{t}^{1},\mathcal{L}(X_{t}^{1}),Y_{t}^{1},Z_{t}^{1})-\bar{b}(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})$ and $\Delta\bar{\sigma}(t):=\bar{\sigma}(t,X_{t}^{1},\mathcal{L}(X_{t}^{1}),Y_{t}^{1},Z_{t}^{1})-\bar{\sigma}(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})$. Taking expectations on both sides and noticing $\bar{Y}_{T}=0$, we get
@@
$\qquad~+\left.(\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{1}),Y_{t}^{2},Z_{t}^{2})-\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2}))(X_{t}^{1}-X_{t}^{2},Y_{t}^{1}-Y_{t}^{2},Z_{t}^{1}-Z_{t}^{2})\right\}\mathrm{d}t$
$\qquad~+\left.\int_{0}^{T}\bar{\eta}_{t}\left\{a(t)\bar{X}_{t}+b(t)\mathbb{E}[\bar{X}_{t}]\right\}\mathrm{d}t+\int_{0}^{T}G(t)\bar{Y}_{t}\mathrm{d}\bar{\zeta}_{t}\right].$
@@
where we used the fact that $\mathbb{E}[\int_{0}^{T}\phi(t,X_{t}^{i},\mathcal{L}(X_{t}^{i}))\eta_{t}^{i}\mathrm{d}t]=0$, $\phi_{i}(t,X_{t}^{i},\mathcal{L}(X_{t}^{i}))\geq 0$ and $\eta_{t}^{i}\geq 0$ for $i=1,2$. Similarly, we have $\mathbb{E}[\int_{0}^{T}G(t)\bar{Y}_{t}\mathrm{d}\bar{\zeta}_{t}]\leq 0$. Exploiting Assumption 3, it holds that
$0$ $\leq\mathbb{E}\Big[\left(\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{1}),Y_{t}^{2},Z_{t}^{2})-\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})\right)\left(\bar{X}_{t},\bar{Y}_{t},\bar{Z}_{t}\right)$
$\qquad\quad-\beta\left(|\bar{X}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2}\right)\Big].$
However, it follows from (A3) in Assumption 1 that

### v3

@@ 4 SMP by the Lagrange Multipliers Method
$\qquad-\int_{0}^{t}\mathbb{E}^{\prime}\left[\partial_{\mu}b(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})(\widehat{X}_{s}^{\prime})^{\top}(X_{s}^{\prime}-\widehat{X}_{s}^{\prime})\right]\mathrm{d}s-\int_{0}^{t}\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})(\widehat{X}_{s}^{\prime})^{\top}(X_{s}^{\prime}-\widehat{X}_{s}^{\prime})\right]\mathrm{d}W_{s}.$
By setting $L=X-\widehat{X}$ and allowing $L$ to vary arbitrarily in $\mathbb{M}^{2}$, we conclude that $\langle\lambda,D_{X}F(\widehat{X},\widehat{\alpha})(L)\rangle_{\mathtt{H}^{*},\mathtt{H}}=0$ for all $L\in\mathbb{M}^{2}$. In light of Lemma 3.9 and Lemma 3.10 of [7], we deduce that $\lambda=0$ that contradicts with the non-degeneracy of the multipliers. Hence, the claim holds and we can scale the multipliers such that (ii) holds. Then, we can conclude from the structure of Lions derivative (c.f. Section 5.2.1 in [11]) and Lemma 4.1 that
$0$ $\leq{\mathbb{E}\left[\left\langle\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T}),L_{T}\right\rangle+\mathbb{E}^{\prime}\left[\left\langle\partial_{\mu}g(\widehat{X}_{T},\widehat{\mu}_{T})(\widehat{X}_{T}^{\prime}),L_{T}^{\prime}\right\rangle\right]\right]}$
$\quad+r_{0}\mathbb{E}\left[\int_{0}^{T}\langle\nabla_{x}f(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t}),L_{t}\rangle\mathrm{d}t+\mathbb{E}^{\prime}\left[\int_{0}^{T}\langle\partial_{\mu}f(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime}),L_{t}^{\prime}\rangle\mathrm{d}t\right]\right]$
$\quad-\sum_{i\in I}\mathbb{E}\left[\int_{0}^{T}\left\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})+\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime})\right],L_{t}^{\prime}\right\rangle\eta^{i}_{t}\mathrm{d}t\right]$
$\quad-\mathbb{E}\left[\langle\lambda_{1}(0),L_{0}\rangle+\int_{0}^{T}\langle\lambda_{1}^{1}(t),L_{t}^{1}-\nabla_{x}b(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})L_{t}\rangle\mathrm{d}t\right$.
$\quad+\int_{0}^{T}\mathrm{tr}\left(\lambda_{1}^{2}(t)(L_{t}^{2}-\nabla_{x}\sigma(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})L_{t})^{\top}\right)\mathrm{d}t$
$\quad+\int_{0}^{T}\left\langle\lambda_{1}^{1}(t),\mathbb{E}^{\prime}\left[\partial_{\mu}b(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime})L_{t}^{\prime}\right]\right\rangle\mathrm{d}t$
$\quad+\left.\int_{0}^{T}\mathrm{tr}\left(\lambda_{1}^{2}(t)\left(\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime})L_{t}^{\prime}\right]\right)^{\top}\right)\mathrm{d}t\right]$
$\quad+\left\langle\lambda_{2},\nu\right\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathcal{M}[0,T])^{n}},$ (26)
where $\lambda_{1}\in\mathbb{M}^{2}$ has the decomposition $\lambda_{1}(\cdot)=\lambda_{1}(0)+\int_{0}^{\cdot}\lambda_{1}^{1}(s)\mathrm{d}s+\int_{0}^{\cdot}\lambda_{1}^{2}(s)\mathrm{d}W_{s}$, $L=X-\widehat{X}$, with $X\in\mathscr{X}$ has the representation $L_{\cdot}=L_{0}+\int_{0}^{\cdot}L_{s}^{1}\mathrm{d}s+\int_{0}^{\cdot}L_{s}^{2}\mathrm{d}W_{s}+\int_{0}^{\cdot}G(s)\mathrm{d}L_{s}^{3}$, and $\nu$ is the associated radon measure of the mapping $\int_{0}^{\cdot}G(s)\mathrm{d}L_{s}^{3}\in\mathrm{BV}^{n}[0,T]$. Define an element $p=(p_{t})_{t\in[0,T]}$ in $\mathbb{M}^{2}$ that, for $t\in[0,T]$,
$p_{t}$ $={r_{0}\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T})+r_{0}\mathbb{E}^{\prime}\left[\partial_{\mu}g(\widehat{X}_{T}^{\prime},\widehat{\mu}_{T})(\widehat{X}_{T})\right]}$
$\quad-\sum_{i\in I}\int_{t}^{T}\left(\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})+\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s}){\eta^{i}_{s}}^{\prime}\right]\right)\mathrm{d}s$
@@
where $MN:=\sum_{i=1}^{r}M_{i}L_{i}\in\mathbb{R}^{n}$ for any $M=(M_{1},\ldots,M_{r})\in\mathbb{R}^{n\times n\times r}$ and $L=(L_{1},\ldots,L_{r})\in\mathbb{R}^{n\times r}$. It then holds that
$\langle\lambda_{1}^{1}(t),L_{t}^{1}\rangle\mathrm{d}t$ $=\langle\lambda_{1}^{1}(t)-p(t),\mathrm{d}L_{t}\rangle+\langle p(t),\mathrm{d}L_{t}\rangle-\langle\lambda^{1}_{1}(t),L^{2}_{t}\mathrm{d}W_{t}\rangle$
$\quad-\langle\lambda_{1}^{1},\nu\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathcal{M}[0,T])^{n}}.$ (28)
@@
$\mathbb{E}\left[\int_{0}^{T}\langle p_{t},\mathrm{d}L_{t}\rangle\right]=-\mathbb{E}\left[\int_{0}^{T}\langle L_{t},\mathrm{d}p_{t}\rangle\right]-\mathbb{E}\left[\langle p,L\rangle_{T}\right]{+\mathbb{E}\left[\langle p_{T},L_{T}\rangle\right]-\mathbb{E}\left[\langle p_{0},L_{0}\rangle\right]}$
@@
$\mathbb{E}\left[\langle p_{0}-\lambda_{1}(0),L_{0}\rangle+\int_{0}^{T}\langle p_{t}-\lambda_{1}^{1}(t),\mathrm{d}L_{t}\rangle+\langle\lambda_{2}-\lambda_{1}^{1},\nu\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathcal{M}[0,T])^{n}}\right]\geq 0.$
By setting $\widehat{X}^{3}=X^{3}$ and $L_{t}=-(p_{0}-\lambda_{1}(0))-\int_{0}^{t}(p_{s}-\lambda_{1}^{1}(s))\mathrm{d}s$, we deduce that $p_{0}=\lambda_{1}(0)$ and $p$ is indistinguishable from $\lambda_{1}^{1}$ by the arbitrariness of $L$. As a result, $\lambda_{1}^{1}$ solves (5.3). Moreover, by letting $X^{1}=\widehat{X}^{1}$ and $\widehat{X}^{2}=X^{2}$, we have $G^{\top}\lambda_{2}=G^{\top}\lambda_{1}^{1}$ in $(C([0,T];\mathbb{R})^{**})^{k}$, and hence $G^{\top}\lambda_{2}\in C([0,T];\mathbb{R}^{k})$. Here, $G$ can be viewed a mapping in $\mathcal{L}((C([0,T];\mathbb{R})^{**})^{n};$ $(C([0,T];\mathbb{R})^{**})^{k})$. By the optimality condition for $\widehat{\alpha}\in\mathscr{U}[0,T]$, we have
$0$ $\leq\mathbb{E}\left[\int_{0}^{T}\left\langle r_{0}\nabla_{u}f(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})+\nabla_{u}b(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})^{\top}\lambda_{1}^{1}(t)+\nabla_{u}\sigma(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})^{\top}\lambda_{1}^{2}(t),\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t\right$.
@@
$0$ $\leq-\mathbb{E}\left[\langle\xi_{3},\eta-\widehat{\zeta}\rangle_{(C([0,T];\mathbb{R})^{**})^{n},({\cal M}[0,T])^{n}}\right]$
$=\mathbb{E}\left[r_{0}\int_{0}^{T}c(t)\mathrm{d}(\eta_{t}-\hat{\zeta}_{t})+\left\langle\lambda_{2},\int_{0}^{T}G(t)\mathrm{d}(\eta_{t}-\hat{\zeta}_{t})\right\rangle_{(C([0,T];\mathbb{R})^{**})^{n},({\cal M}[0,T])^{n}}\right]$
@@ 5 Uniqueness and Stability of Solution to Constrained FBSDE
$Y_{t}$ $={r_{0}\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T})+r_{0}\mathbb{E}^{\prime}\left[\partial_{\mu}g(\widehat{X}_{T}^{\prime},\widehat{\mu}_{T})(\widehat{X}_{T})\right]}$
$\quad+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\mu},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H^{r_{0}}(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime},Y_{s}^{\prime},Z_{s}^{\prime})(\widehat{X}_{s})\right]\mathrm{d}s$
@@
subject to $|r_{0}|+\|\eta\|_{\mathbb{L}_{\mathbb{F},\mathbb{R}}}=1$, $\widehat{\mu}_{t}=\mathcal{L}(\widehat{X}_{t})$, $\mathbb{P}\left(G(t)Y_{t}+{r_{0}}c(t)\geq 0,~\forall t\in[0,T]\right)=1$, and $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.
@@
We now verify the constraint qualification (CQ) condition as in (1.4) of [38] to establish the Karush-Kuhn-Tucker (KKT) optimality condition (i.e., we verify $r_{0}=1$ in Lemma 4.1).
@@
Here, the Hamiltonian $H(t,x,\mu,u,y,z):=H^{1}(t,x,\mu,u,y,z)$ for $(t,x,\mu,u,y,z)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$, and the process pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ is the unique solution to the following general type of constrained BSDE:
$Y_{t}$ $={\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T})+\mathbb{E}^{\prime}\left[\partial_{\mu}g(\widehat{X}_{T}^{\prime},\widehat{\mu}_{T})(\widehat{X}_{T})\right]}$
$\quad+\int_{t}^{T}\nabla_{x}{H}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime},Y_{s}^{\prime},Z_{s}^{\prime})(\widehat{X}_{s})\right]\mathrm{d}s$
@@
As a result, we can set $r_{0}=1$ in (5)-(31) under Assumption 3. The strict convexity imposed on the Hamiltonian $H$ w.r.t. the control variable $u\in U$ given in (C3) of Assumption 3 yields the result that once $\langle\nabla_{u}H(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t},Y_{t},Z_{t}),u-\widehat{\alpha}_{t}\rangle\geq 0$ for all $u\in U$, we must have $\widehat{\alpha}_{t}=\hat{u}(t,\widehat{X}_{t},\widehat{\mu}_{t},Y_{t},Z_{t})$. Then, we have
@@
Let $(X^{i}_{t},\zeta^{i}_{t},Y^{i}_{t},Z^{i}_{t},\eta_{t}^{i})_{t\in[0,T]}$ ($i=1,2$) be two solutions to the FBSDE (5)-(31) satisfying $X^{1}_{0}=X_{0}^{2}$ and $Y_{T}^{1}=Y_{T}^{2}$. It follows from Lemma 5.3 that the KKT condition holds under Assumption 2, and hence we can set the coefficient $r_{0}=1$ in the Hamiltonian $H^{r_{0}}$. Recall $F(t,x,\mu,u,y,z):=\nabla_{x}H(t,x,\mu,u,y,z)+\int_{\mathbb{R}^{n}}\partial_{\mu}H(t,x^{\prime},\mu,u,y,z)(x)\mu(\mathrm{d}x^{\prime})$ defined in Assumption 3-(C3). Let us set $\bar{X}_{t}:=X_{t}^{1}-X_{t}^{2}$, $\bar{\zeta}_{t}:=\zeta_{t}^{1}-\zeta_{t}^{2}$, $\bar{Y}_{t}:=Y_{t}^{1}-Y_{t}^{2}$, $\bar{Z}_{t}:=Z_{t}^{1}-Z_{t}^{2}$ and $\bar{\eta}_{t}:=\eta_{t}^{1}-\eta_{t}^{2}$ for $t\in[0,T]$. Applying Itô’s formula to $\bar{X}_{t}\bar{Y}_{t}$, we arrive at
$\bar{X}_{T}\bar{Y}_{T}$ $=\int_{0}^{T}\left\{-\Delta F(t)\bar{X}_{t}+\bar{Y}\Delta{b}(t)+\mathrm{tr}(\bar{Z}_{t}\Delta{\sigma}(t))\right\}\mathrm{d}t+\int_{0}^{T}\left(\bar{X}_{t}\bar{Z}_{t}+\bar{Y}_{t}\Delta\sigma(t)\right)\mathrm{d}W_{t}$
@@
where the coefficients $\Delta\Psi=\Psi(X_{T}^{1},\mathcal{L}(X_{T}^{1}))-\Psi(X_{T}^{2},\mathcal{L}(X_{T}^{2}))$, $\Delta F(t):=F(t,X_{t}^{1},\mathcal{L}(X_{t}^{1}),Y_{t}^{1},Z_{t}^{1})-F(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})$, $\Delta b(t):=b(t,X_{t}^{1},\mathcal{L}(X_{t}^{1}),Y_{t}^{1},Z_{t}^{1})-b(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})$ and $\Delta\sigma(t):=\sigma(t,X_{t}^{1},\mathcal{L}(X_{t}^{1}),Y_{t}^{1},Z_{t}^{1})-\sigma(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})$
Taking expectations on both sides and noticing that $\bar{Y}_{T}=\Delta\Psi$, we get
@@
$\hskip 18.49988pt~+\left.(\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{1}),Y_{t}^{2},Z_{t}^{2})-\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2}))(X_{t}^{1}-X_{t}^{2},Y_{t}^{1}-Y_{t}^{2},Z_{t}^{1}-Z_{t}^{2})\right\}\mathrm{d}t$
$\hskip 18.49988pt~+\left.\int_{0}^{T}\bar{\eta}_{t}\left\{a(t)\bar{X}_{t}+b(t)\mathbb{E}[\bar{X}_{t}]\right\}\mathrm{d}t+\int_{0}^{T}G(t)\bar{Y}_{t}\mathrm{d}\bar{\zeta}_{t}\right].$
@@
where, we used the fact that $\mathbb{E}[\int_{0}^{T}\phi(t,X_{t}^{i},\mathcal{L}(X_{t}^{i}))\eta_{t}^{i}\mathrm{d}t]=0$, $\phi_{i}(t,X_{t}^{i},\mathcal{L}(X_{t}^{i}))\geq 0$ and $\eta_{t}^{i}\geq 0$ for $i=1,2$. Similarly, we have $\mathbb{E}[\int_{0}^{T}G(t)\bar{Y}_{t}\mathrm{d}\bar{\zeta}_{t}]\leq 0$. Exploiting Assumption 3, we deduce that
$0$ $\leq\mathbb{E}\left[\bar{X}_{T}\Delta\Psi\right]\leq\mathbb{E}\big[\left(\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{1}),Y_{t}^{2},Z_{t}^{2})-\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})\right)\left(\bar{X}_{t},\bar{Y}_{t},\bar{Z}_{t}\right)$
$\qquad\qquad\qquad\qquad\quad-\beta\left(|\bar{X}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2}\right)\big].$
However, it follows from Assumption 1-(A3) that